OpenStudy (anonymous):

Divide and state any restrictions on the variables. (x^2-16)/(x^2+5x+6)÷(x^2+5x+4)/(x^2-2x-8)

OpenStudy (anonymous):

\[{\frac{x^2-16}{x^2+5x+6}}\div \frac{x^2+5x+4}{x^2-2x-8}\] Let's first factor the top and the bottom. That's almost always the first step in these types of problems. \[\frac{(x+4)(x-4)}{(x+3)(x+2)}\div \frac{(x+4)(x+1)}{(x-4)(x+2)}\] Let's change this to a multiplication problem. \[\frac{(x+4)(x-4)}{(x+3)(x+2)}\frac{(x-4)(x+2)}{(x+4)(x+1)}\] Let's start cancelling terms that appear on the top and the bottom. (x+4) and (x+2) will cancel. \[\frac{(x-4)^2}{(x+3)(x+1)}\]\[x \neq -1,-2,-3, or-4\]